OpenStudy (anonymous):

The numerator of a certain fraction is 5 less than the denominator. If the fraction is equal to 3/4, what is the denominator of this fraction?

OpenStudy (mertsj):

Let d = the denominator the numerator = d-5 So:\[\frac{d-5}{d}=\frac{3}{4}\]

OpenStudy (mertsj):

Cross multiply and solve. Can you do that?

Directrix (directrix):

4(d - 5) = 3d 4d - 20 = 3d d = 20